Write the slope-intercept equation for the line that passes through (4, 3) and has slope 5/4. Please show all of your work.

Answer by mananth(16946)

Equation in slope intercept form
y=mx+b
slope =5/4
point =(4,3)
Find b
plug the values
3=5/4*4 +b
3=5+b
b=-2
..
slope = 5/4 b=-2
equation is
y= 5x/4 -2